OpenStudy (anonymous):

A baby elephant weighs 200 pounds at birth. Seven years later, the elephant's weight is 3,000 pounds. What is the rate of change of the weight of the elephant? A) 400 pounds per year B) 429 pounds per year C) 457 pounds per year D) 466 pounds per year

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

the rate begins at 200. therefore you need a 3000-200 which 2800 lbs gained over 7 years. 400 lbs per year

OpenStudy (anonymous):

ok thank you :)

OpenStudy (bradely):

Change in weight = 3000-200 = 2800 Change in time = 7-0 = 7 Rate of change = Change in weight /Change in time =2800/7 =400 A is the answer Source: http://www.mathskey.com/question2answer/